Brent Bush, CFO of a medical device manufacturer, BioTron Medical, Inc., was approached by a Japanese customer, Numata, with a proposal to pay cash (in yen) for its typical orders of ¥12,500,000 every other month if it were given a 4.5% discount. Numata's current terms are 30 days with no discounts. Using the following quotes and estimated cost of capital for Numata, Bush will compare the proposal with covering yen payments with forward contracts. What is the difference in US$ between the payments made at discount and no discount, but fully covered with a forward contract? Given the following information:

Assumptions Values
BioTron's 30-day account receivable, Japanese yen 10823614
Spot rate, ¥/$                   111.93
30-day forward rate, ¥/$                   111.09
90-day forward rate, ¥/$                   109.60
180-day forward rate, ¥/$                   109.20
Numata's WACC 8.850%
BioTron Medical's WACC 9.49%
Desired discount on purchase price by Numata 4.500%
Answer #1

1. Allow the discount and receive payment in Japanese yen

Account received in yen= 12,500,000

Discount for cash payment is 4.500%=12,500,00 *0.045=5,62,500 yen

Amount paid after deducting discount= 12500000 - 562500=11,937,500 yen

Current spot rate ,yen/dollar= 111.40

Amount received in U.S. dollar =11,937,500/11.40= $107,158.89

2.Not offer any discount for early payment and cover exposer with forward

Account receivable (yen) = 12,500,000

30 days forward rate = 111.09

Amount received in cash in dollar in 30 days =12,500,000/111.09 = $112,521.38

Difference in the amount without discount and with discount=$112,521.38 - $107,158.89= $5363 answer.

Brent Bush should politely decline Numata's offer to pay cash in exchange for the requested discount.